### v1

# The Bohlin variant of the Eisenhart lift

The Bohlin variant of the Eisenhart lift
Anton Galajinsky Tomsk Polytechnic University, 634050 Tomsk, Lenin Ave. 30, Russia
Tomsk State University of Control Systems and Radioelectronics,
Lenin ave. 40, 634050 Tomsk, Russia
e-mail: galajin@tpu.ru
Inspired by the Bohlin transformation relating the planar harmonic oscillator to the Kepler problem, a variant of the Eisenhart lift is studied, in which a Lagrangian conservative dynamical system with $d$ degrees of freedom is embedded into timelike geodesics of a conformally flat metric on a (d+2)(d+2)–dimensional space–time of the Lorentzian signature. The uplift is used to construct novel examples of conformally flat metrics admitting higher rank Killing tensors.
Keywords: Eisenhart lift, Bohlin transformation, conformally flat metrics, hidden symmetries
1. Introduction
Since the formulation of general relativity, geometry and theoretical physics go parallel and mutually enriching courses. In particular, geometrization of a Lagrangian conservative mechanics amounts to embedding Newton’s equation into the null geodesics of the Eisenhart metric [1, 2]
$L(x,\dot{x})dt^{2}+dtds,$
where $L(x,\dot{x})$ is the Lagrangian density, $t$ is the temporal coordinate, and $s$ is an extra variable [1, 2]. The latter is needed in order to reproduce the conventional nonrelativistic energy conservation condition within the relativistic framework. At the level of equations of motion, $s$ links to the action variable sometimes adopted in classical mechanics. Thus, given a dynamical system with $d$ degrees of freedom $x_{i}$, $i=1,\dots,d$, the resulting space–time is (d+2)(d+2)–dimensional and the metric is of the Lorentzian signature.
A salient feature of Eisenhart’s metric is that it admits a covariantly constant null Killing vector field $\frac{\partial}{\partial s}$, meaning that the space–time belongs to the Kundt class, as it contains a geodesic null congruence with vanishing expansion, shear and vorticity.
If a Lagrangian density explicitly depends on time, $L(t,x,\dot{x})$, the metric is referred to as the Brinkmann–type metric or a plane–fronted wave with parallel rays (the pppp–wave).
As is well known, in Riemannian geometry symmetries of a metric are described by Killing vectors, which generate infinitesimal coordinate transformations in space–time, $y^{\prime A}=y^{A}+\xi^{A}(y)$. As far as the geodesic equations are concerned, each Killing vector field with components $\xi^{A}(y)$ gives rise to the first integral $\xi^{A}(y)g_{AB}(y)\frac{dy^{A}}{d\tau}$ of the geodesic equations, where $g_{AB}(y)$ is a metric and $\tau$ is the proper time.
A space–time may also admit Killing tensors, i.e. totally symmetric tensor fields obeying the equation $\nabla_{(A_{1}}K_{A_{2}\dots A_{n+1})}(y)=0$, which underlie first integrals of the form $K_{A_{1}\dots A_{n}}(y)\frac{dy^{A_{1}}}{d\tau}$
$\dots\frac{dy^{A_{n}}}{d\tau}$. The existence of Killing tensors is usually attributed to hidden symmetries of a space–time, as there is no coordinate transformation associated to them.
The importance of the Killing vectors and tensors is hard to overestimate. They give a clue to establishing complete integrability of the geodesic equations formulated in various black hole spacetimes, as well as allow one to separate variables in the Hamilton–Jacobi, Klein–Gordon and Dirac equations in strong gravitational fields (for a review and further references see [3]).
Apart from the aesthetic appeal, the Eisenhart lift provides an efficient means of constructing novel space–times with hidden symmetries. By applying the method to various integrable models one can build a plethora of examples of such a kind [4][16]. Some of them provide exact solutions to the Einstein equations [4, 11, 12, 15]. Worth mentioning also is a considerable body of recent work on application of the Eisenhart lift to cosmology (see e.g. [17][20] and references therein).
The goal of this work is to study a variant of the Eisenhart lift, which is inspired by the Bohlin transformation [21] (for a review see [22]) connecting the planar harmonic oscillator to the Kepler problem. Given a dynamical system with $d$ degrees of freedom, a simple conformally flat metric on a (d+2)(d+2)–dimensional space–time of the Lorentzian signature is proposed, timelike geodesics of which include the original Newton equation.
The work is organized as follows.
In the next section, we briefly outline the Eisenhart lift, paying particular attention to physical dimensions of the ingredients involved.
In Sect. 3, by formally applying the Bohlin transformation to the Eisenhart metric originating from the two–dimensional harmonic oscillator, we obtain a simple conformally flat metric (the conformal factor being the Kepler potential), which can be readily generalized to the case of arbitrary dimension and potential. In contrast to the Eisenhart framework, the metric does not belong to the Kundt class and the original Newton equation is recovered by analyzing timelike geodesics rather than null geodesics. The issue of hidden symmetries is discussed as well.
Sect. 4 contains explicit examples. Firstly, it is shown that the celebrated conformal mechanics of [23], when Bohlin uplifted, yields a metric of the three–dimensional anti de Sitter space. Secondly, the four–body Calogero model [24] is used to build a six–dimensional conformally flat metric admitting irreducible Killing tensors of rank 3 and 4. By increasing the number of particles in the original Calogero model, one can build a (d+2)(d+2)–dimensional Bohlin–type metric admitting Killing tensors of rank up to $d$.
We summarize our results and discuss possible further developments in the concluding Sect. 5.
Throughout the paper, summation over repeated indices is understood unless otherwise stated.
2. The Eisenhart lift
Consider a conservative dynamical system with $d$ degrees of freedom $x_{i}$, $i=1,\dots,d$, and let $t$ be the temporal variable. Assuming that potential energy $V(x)$ can be factorized $V(x)=gU(x)$, where $g$ is a constant having the dimension of energy and $U(x)$ is a positive dimensionless function, [^1] Because the coordinates $x_{i}$ have the dimension of length and $U(x)$ is assumed to be dimensionless, the latter will necessarily involve a (coupling) constant with the dimension of length. By this reason, a numerical value of $g$ entering the original potential $V(x)=gU(x)$ is of no great importance. the Newton equation
$m\frac{d^{2}x_{i}}{dt^{2}}+\frac{\partial V(x)}{\partial x_{i}}=0,$ (1)
where $m$ is particle’s mass, can be embedded into the null geodesics
$\frac{d^{2}y^{A}}{d\lambda^{2}}+\Gamma^{A}_{BC}\frac{dy^{B}}{d\lambda}\frac{dy^{C}}{d\lambda}=0,\qquad g_{AB}\frac{dy^{A}}{d\lambda}\frac{dy^{B}}{d\lambda}=0$ (2)
of the Eisenhart metric [1, 2]
$g_{AB}dy^{A}dy^{B}=2U(x)c^{2}dt^{2}-2cdtds-dx_{i}dx_{i},$ (3)
where $U(x)$ is the prepotential above, $s$ is an extra coordinate having the dimension of length, $c$ is the speed of light, $y^{A}=(ct,s,x_{1},\dots,x_{d})$, and $\lambda$ is an affine parameter (of the dimension of time) along a null geodesic.
Indeed, computing the Christoffel symbols
$\Gamma^{i}_{tt}=-\Gamma^{s}_{ti}=\partial_{i}U,$ (4)
where the notation $A=(t,s,i)$, $\partial_{i}=\frac{\partial}{\partial x_{i}}$, with $i=1,\dots,d$, was adopted, one can reduce the first equation in (2) to
$\frac{dt}{d\lambda}=\alpha,\qquad\frac{ds}{d\lambda}-2\alpha cU=c\beta,\qquad m\frac{d^{2}x_{i}}{dt^{2}}+mc^{2}\partial_{i}U=0,$ (5)
where $\alpha>0$ (forward in time motion) and $\beta$ are dimensionless real constants of integration. Note that in the last equation $\frac{d}{d\lambda}=\frac{dt}{d\lambda}\frac{d}{dt}$ was used and the common factor ${\left(\frac{dt}{d\lambda}\right)}^{2}$ was dropped.
The second equation in (2) relates the ratio $\frac{\beta}{\alpha}$ to particle’s energy
$\frac{m}{2}\frac{dx_{i}}{dt}\frac{dx_{i}}{dt}+mc^{2}U(x)=-\frac{mc^{2}\beta}{\alpha},$ (6)
as well as allows one to determine how the extra variable $s$ evolves with time
$s(t)=-\frac{1}{mc}\int_{0}^{t}d\tilde{t}\left(\frac{m}{2}\frac{dx_{i}\left(\tilde{t}\right)}{d\tilde{t}}\frac{dx_{i}\left(\tilde{t}\right)}{d\tilde{t}}-mc^{2}U\left(x\left(\tilde{t}\right)\right)\right)+s_{0},$ (7)
where $s_{0}$ is a constant of integration. The latter formula implies that $s$ links to the action variable sometimes adopted in classical mechanics. [^2] In principle, one is allowed to consider timelike geodesics within the Eisenhart framework. In that case, the right hand side of (7) would involve an extra contribution $\frac{ct}{2\alpha^{2}}$, which would only spoil the interpretation of $s(t)$ as the action variable. A numerical value of the energy in (6) would also include $\frac{mc^{2}}{2\alpha^{2}}$, thus yielding $\frac{mc^{2}(1+2\alpha\beta)}{2\alpha^{2}}$ on the right hand side of (6).
The Newton equation (1) is thus recovered by identifying $mc^{2}U(x)=V(x)$ and implementing the null reduction along $s$. Note that in physics literature it is customary to use units in which $c=1$ and the mass of a particle is usually set to unity.
Let us briefly discuss properties of the Eisenhart metric (3). Computing the eigenvalues of $g_{AB}$
$U+\sqrt{U^{2}+1},\qquad U-\sqrt{U^{2}+1},\qquad-1,\qquad\dots\qquad,-1,$ (8)
one concludes that (3) describes a (d+2)(d+2)–dimensional space–time of the Lorentzian signature.
The metric admits two Killing vector fields
$\xi^{A}\frac{\partial}{\partial y^{A}}=\frac{\partial}{\partial t},\qquad\eta^{A}\frac{\partial}{\partial y^{A}}=\frac{\partial}{\partial s}.$ (9)
The corresponding first integrals of the geodesic equations
$\xi^{A}g_{AB}\frac{dy^{B}}{d\lambda},\qquad\eta^{A}g_{AB}\frac{dy^{B}}{d\lambda}$
link to the first two equations in (5). In general relativity, a space–time is called stationary if it admits a timelike Killing vector field. For the case at hand, $\xi^{A}$ is timelike provided the prepotential is a positive function $U(x)>0$. The second Killing vector $\eta^{A}$ is null and, more importantly, it is covariantly constant $\nabla_{A}\eta_{B}=0$. Hence, (3) admits a geodesic null congruence with vanishing expansion, shear and vorticity and thus belongs to the Kundt class.
Computing the Ricci tensor, one reveals the only non–vanishing component
$R_{tt}=\Delta U,$
which also means that the scalar curvature vanishes, harmonic prepotentials uplift to Ricci–flat metrics, and no cosmological term is allowed.
Finally, assuming the original dynamical system (1) admits a polynomial integral of motion
$\mathcal{I}=\sum_{k=0}^{n}I_{i_{1}\dots i_{k}}(x)\frac{dx_{i_{1}}}{dt}\dots\frac{dx_{i_{k}}}{dt},\qquad\frac{d\mathcal{I}}{dt}=0,$ (10)
one can uplift it to a Killing tensor [^3] For a discussion of conformal Killing tensors within the Eisenhart framework see e.g. [7, 10]. of rank $n$ just by multiplying $\mathcal{I}$ with ${\left(\frac{dt}{d\lambda}\right)}^{n}$ and representing the latter in the form
$K_{A_{1}\dots A_{n}}(y)\frac{dy^{A_{1}}}{d\lambda}\dots\frac{dy^{A_{n}}}{d\lambda}.$
The resulting expression will automatically obey $\nabla_{(A_{1}}K_{A_{2}\dots A_{n+1})}=0$.
As was mentioned in the Introduction, within the context of general relativity Killing tensors describe hidden symmetries of a space–time. The Eisenhart lift thus provides a convenient framework for building metrics admitting higher rank Killing tensors [4][15], some of which are Ricci–flat [4, 11, 12, 15].
3. The Bohlin variant of the Eisenhart lift
The Bohlin transformation [21] (for a review see [22]) links the planar harmonic oscillator described by a complex coordinate $w$ and the temporal variable $t$ to the Kepler problem specified by a complex coordinate $z$, the evolution parameter $\lambda$, and a negative value of conserved energy
$z=w^{2},\qquad w\bar{w}\frac{dt}{d\lambda}=1.$ (11)
Note that it is the potential of the original dynamical system (evaluated on particle’s orbit) which determines the new temporal coordinate $\lambda$. For simplicity of the presentation, in (11) we ignored physical dimensions.
Building the Eisenhart metric associated with the two–dimensional harmonic oscillator and formally [^4] The second equation in (11) involves the general solution $w(t)$ to the oscillator equation of motion, meaning that (11) cannot be regarded as a proper coordinate transformation within the context of general relativity. implementing (11) there, one obtains a simple conformally flat metric, the conformal factor being the Kepler potential, which can be readily generalized to the case of arbitrary dimension and prepotential
$g_{AB}(y)dy^{A}dy^{B}=U(x)\left(2cdtds-dx_{i}dx_{i}\right),$ (12)
where $y^{A}=(ct,s,x_{1},\dots,x_{d})$, $c$ is the speed of light, and $U(x)$ is an arbitrary positive function.
Let us discuss properties of (12). The eigenvalues of $g_{AB}$ are $\left(U,-U,\dots,-U\right)$ and, hence, it describes a (d+2)(d+2)–dimensional conformally flat stationary space–time of the Lorentzian signature.
For an arbitrary $U(x)$ it admits three Killing vector fields
$\frac{\partial}{\partial t},\qquad\frac{\partial}{\partial s},\qquad t\frac{\partial}{\partial t}-s\frac{\partial}{\partial s},$ (13)
which form Lie algebra of the Poincaré group in $1+1$ dimensions. None of them is covariantly constant and, hence, the space–time does not belong to the Kundt class. If $U(x)$ is a homogeneous function of degree 2-2, the metric holds invariant under the dilatation transformation $y^{\prime A}=\sigma y^{A}$, where $\sigma$ is a constant. Such metrics might be of potential interest for holographic applications.
Computing the Christoffel symbols
$\Gamma^{t}_{ti}=\Gamma^{s}_{si}=\Gamma^{i}_{ts}=\frac{1}{2}U^{-1}\partial_{i}U,\qquad\Gamma^{i}_{jk}=\frac{1}{2}U^{-1}\left(\partial_{j}U\delta_{ik}+\partial_{k}U\delta_{ij}-\partial_{i}U\delta_{jk}\right),$
$\Gamma^{A}_{Ai}=\frac{(d+2)}{2}U^{-1}\partial_{i}U,$ (14)
where $\delta_{ij}$ is the Kronecker delta and $\partial_{i}=\frac{\partial}{\partial x_{i}}$, one can then determine the Ricci tensor
$R_{ts}=\frac{1}{2}U^{-1}\left(\Delta U+\frac{(d-2)}{2}U^{-1}\partial_{i}U\partial_{i}U\right),$
$R_{ij}=-\frac{d}{2}U^{-1}\left(\partial_{i}\partial_{j}U-\frac{3}{2}U^{-1}\partial_{i}U\partial_{j}U\right)-R_{ts}\delta_{ij}$ (15)
and the scalar curvature
$R=(d+1)U^{-2}\left(\Delta U+\frac{(d-4)}{4}U^{-1}\partial_{i}U\partial_{i}U\right).$ (16)
Taking into account the identity
$\partial_{i}\partial_{j}U^{k}=kU^{k-1}\left(\partial_{i}\partial_{j}U+(k-1)U^{-1}\partial_{i}U\partial_{j}U\right),$ (17)
one concludes that the space–time is Ricci–flat only for a trivial prepotential $U=\mbox{const}$, while the scalar curvature vanishes for
$U=F^{\frac{4}{d}},\qquad\Delta F=0.$ (18)
Now let us turn to the timelike geodesics
$\frac{d^{2}y^{A}}{d\tau^{2}}+\Gamma^{A}_{BC}\frac{dy^{B}}{d\tau}\frac{dy^{C}}{d\tau}=0,\qquad g_{AB}\frac{dy^{A}}{d\tau}\frac{dy^{B}}{d\tau}=c^{2},$ (19)
where $\tau$ is the proper time, which follow from (12). Taking into account (S0.Ex5), the equations (19) can be reduced to
$U\frac{dt}{d\tau}=\alpha,\qquad\frac{1}{c}U\frac{ds}{d\tau}=\beta,\qquad m\frac{d^{2}x_{i}}{dt^{2}}+\frac{mc^{2}}{2\alpha^{2}}\partial_{i}U=0,$ (20)
where $\alpha>0$ (forward in time motion) and $\beta$ are dimensionless real constants of integration, while the ratio $\frac{\beta}{\alpha}$ links to particle’s energy
$\frac{m}{2}\frac{dx_{i}}{dt}\frac{dx_{i}}{dt}+\frac{mc^{2}}{2\alpha^{2}}U(x)=\frac{mc^{2}\beta}{\alpha}.$ (21)
In contrast to the Eisenhart framework, the dynamics of the extra variable $s$ is affinely related to ctct
$\frac{ds}{dt}=\frac{c\beta}{\alpha}.$ (22)
Identifying $\frac{mc^{2}}{2\alpha^{2}}U(x)=V(x)$ and implementing the null reduction along $s$, one recovers Newton’s equation (1).
Although Bohlin’s work was published eighteen years prior to Eisenhart’s paper, it seems tempting to call (12) the Bohlin variant of the Eisenhart lift, because the coordinate time and the proper time are related by the prepotential similarly to the second equation in (11).
Note that, if (10) happens to be a polynomial integral of motion of a dynamical system at hand, one can turn it into a Killing tensor of the Bohlin metric. The uplift is less straightforward than within the Eisenhart framework, however. Because Newton’s equation (1) holds invariant under the time reversal transformation $t\to-t$, the corresponding first integrals involve monomials in momenta which are all of even/odd power. Substituting $p_{i}=\frac{dx_{i}}{dt}=\frac{U}{\alpha}\frac{dx_{i}}{d\tau}$ into the integrals of motion and multiplying, where necessary, with the unity $1=\frac{1}{c^{2}}g_{AB}\frac{dy^{A}}{d\tau}\frac{dy^{B}}{d\tau}$, one obtains a homogeneous function $K_{A_{1}\dots A_{n}}\frac{dy^{A_{1}}}{d\tau}\dots\frac{dy^{A_{n}}}{d\tau}$, from which components of the Killing tensor are read off.
The Bohlin lift thus can be used to build novel conformally flat metrics of the Lorentzian signature admitting higher rank Killing tensors. Examples of such a kind are given in the next section.
Concluding this section, we note that a possibility to accommodate the Bohlin transformation within the Jacobi geometrization of classical dynamics was studied in [25].
4. Examples
Our first example originates from the prepotential
$U(x_{1},\dots,x_{d})=\frac{1}{{\left(a+b_{i}x_{i}\right)}^{2}},$ (23)
where $a$ is a dimensionless constant and $b_{i}$ are constants having the dimension of inverse length. When Bohlin uplifted, it results in a metric solving the vacuum Einstein equations in the presence of a cosmological constant [^5] As a matter of fact, (23) yields a metric defined on a patch of the (d+2)(d+2)–dimensional anti de Sitter space, because $R_{ABCD}=(b_{i}b_{i})\left(g_{AC}g_{BD}-g_{AD}g_{BC}\right)$.
$R_{AB}-\frac{1}{2}Rg_{AB}=\Lambda g_{AB},$
with $R=(d+1)(d+2)b_{i}b_{i}$ and $\Lambda=-\frac{d(d+1)b_{i}b_{i}}{2}$. In particular, the celebrated conformal mechanics [23], which is governed by the potential $V=\frac{g^{2}}{x^{2}}$, $g$ being a coupling constant, belongs to this class. As far as the metric is concerned, the constant $a$ in (23) can be removed by redefining $x_{i}$. The resulting line element is invariant under the extra dilatation transformation $y^{\prime A}=\sigma y^{A}$, where $\sigma$ is a constant.
Our second example also belongs to the class of scaling invariant metrics and builds upon the Calogero model [24], which describes a set of $d$ identical particles on the real line interacting via the inverse–square potential
$V(x_{1},\dots,x_{d})=\frac{1}{2}\sum_{i\neq j}^{d}\frac{g^{2}}{{(x_{i}-x_{j})}^{2}},$ (24)
where $g$ is a coupling constant. The model is integrable and the corresponding first integrals can be obtained from the Lax matrix (no sum over repeated indices below)
$L_{ij}=p_{i}\delta_{ij}+{\rm i}g\left(1-\delta_{ij}\right){\left(x_{i}-x_{j}\right)}^{-1},$
where $p_{i}(t)=\frac{dx_{i}(t)}{dt}$ are momenta canonically conjugate to coordinates, $\{x_{i},p_{j}\}=\delta_{ij}$, and $\rm i$} is the imaginary unit. Constants of motion are polynomial in the momenta, $I_{i}=\frac{1}{i!}{\rm tr}L^{i}$, with $i=1,\dots,d$. For example, the first four members in the string read
$I_{1}=\sum_{i=1}^{d}p_{i},\quad I_{2}=\frac{1}{2}\sum_{i=1}^{d}p_{i}^{2}+\sum_{i<j}^{d}\frac{g^{2}}{{(x_{i}-x_{j})}^{2}},\quad I_{3}=\frac{1}{3!}\left(\sum_{i=1}^{d}p_{i}^{3}+3g^{2}\sum_{i<j}^{d}\frac{p_{i}+p_{j}}{{(x_{i}-x_{j})}^{2}}\right),$
$I_{4}=\frac{1}{4!}\left(\sum_{i=1}^{d}p_{i}^{4}+4g^{2}\sum_{i<j}^{d}\frac{p_{i}^{2}+p_{j}^{2}+p_{i}p_{j}}{{(x_{i}-x_{j})}^{2}}+2g^{4}\sum_{i<j}^{d}\frac{1}{{(x_{i}-x_{j})}^{4}}+\right$.
$\qquad\qquad\left.+4g^{4}\sum_{i\neq j,i\neq k,j<k}^{d}\frac{1}{{(x_{i}-x_{j})}^{2}{(x_{i}-x_{k})}^{2}}\right).$ (25)
Let us focus on the four–body case and construct a six–dimensional Bohlin–type metric, which admits irreducible Killing tensors of rank 3 and 4. For simplicity of the presentation, below we set $m=c=1$.
The prepotential which specifies the metric in (12) reads
$U(x_{1},\dots,x_{d})=\frac{1}{2}\sum_{i\neq j}^{4}\frac{L^{2}}{{(x_{i}-x_{j})}^{2}},$
where $L$ is a constant having the dimension of length, which links to the coupling constant $g$ in (24) via $g^{2}=\frac{L^{2}}{2\alpha^{2}}$. As above, $\alpha>0$ is a constant entering the leftmost relation in (20).
Substituting $p_{i}=\frac{dx_{i}}{dt}=\frac{U}{\alpha}\frac{dx_{i}}{d\tau}$ into $I_{3}$ and $I_{4}$ in (S0.Ex9) and multiplying the terms involving $g^{2}$ by the unity $1=g_{AB}\frac{dy^{A}}{d\tau}\frac{dy^{B}}{d\tau}$ and those containing $g^{4}$ by $1={\left(g_{AB}\frac{dy^{A}}{d\tau}\frac{dy^{B}}{d\tau}\right)}^{2}$, so as to obtain expressions homogeneous in $\frac{dy^{A}}{d\tau}$, one finally obtains the third–rank Killing tensor (no sum over repeated indices unless otherwise stated)
$K_{tsi}=\frac{1}{2}L^{2}U^{2}\sum_{a,b,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}},$
$K_{ijp}=U^{3}\delta_{ip}\delta_{jp}-\frac{1}{2}L^{2}U^{2}\left(\delta_{ij}\sum_{a\neq p}^{4}\frac{1}{{(x_{a}-x_{p})}^{2}}+\delta_{jp}\sum_{a\neq i}^{4}\frac{1}{{(x_{a}-x_{i})}^{2}}+\delta_{pi}\sum_{a\neq j}^{4}\frac{1}{{(x_{a}-x_{j})}^{2}}\right),$
as well as the fourth–rank Killing tensor
$K_{ttss}=\frac{1}{6}L^{4}U^{2}\left(\sum_{a,b,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right),$
$K_{tsij}=-\frac{1}{12}L^{4}U^{2}\delta_{ij}\left(\sum_{a,b,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right)$
$+\frac{1}{6}L^{2}U^{3}\sum_{a,b,a\neq b}^{4}\frac{\left(2\delta_{ai}\delta_{aj}+\delta_{ai}\delta_{bj}\right)}{{(x_{a}-x_{b})}^{2}},$
$K_{ijkp}=U^{4}\sum_{a=1}^{4}\delta_{ai}\delta_{aj}\delta_{ak}\delta_{ap}$
$+\frac{1}{12}L^{4}U^{2}\left(\delta_{ij}\delta_{kp}+\delta_{ik}\delta_{pj}+\delta_{ip}\delta_{jk}\right)\left(\sum_{a,b,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right)$
$-\frac{1}{6}L^{2}U^{3}\left(2\sum_{a,b,a\neq b}^{4}\frac{\left(\delta_{ai}\delta_{aj}\delta_{kp}+\delta_{ai}\delta_{ak}\delta_{pj}+\delta_{ai}\delta_{ap}\delta_{jk}+\delta_{aj}\delta_{ak}\delta_{pi}+\delta_{aj}\delta_{ap}\delta_{ik}+\delta_{ak}\delta_{ap}\delta_{ij}\right)}{{(x_{a}-x_{b})}^{2}}\right$.
$\left.+\sum_{a,b,a\neq b}^{4}\frac{\left(\delta_{ai}\delta_{bj}\delta_{kp}+\delta_{ai}\delta_{bk}\delta_{pj}+\delta_{ai}\delta_{bp}\delta_{jk}+\delta_{aj}\delta_{bk}\delta_{pi}+\delta_{aj}\delta_{bp}\delta_{ik}+\delta_{ak}\delta_{bp}\delta_{ij}\right)}{{(x_{a}-x_{b})}^{2}}\right).$
In a similar fashion, by increasing the number of particles in the original Calogero model, one can build a (d+2)(d+2)–dimensional Bohlin–type metric admitting Killing tensors of rank up to $d$.
5. Conclusion
To summarize, inspired by the Bohlin transformation relating the planar harmonic oscillator to the Kepler problem, a variant of the Eisenhart lift was proposed, in which a Lagrangian conservative dynamical system with $d$ degrees of freedom was embedded into timelike geodesics of a conformally flat metric defined on a (d+2)(d+2)–dimensional space–time of the Lorentzian signature. The (anti) de Sitter space–time fitted nicely into the framework. The uplift was used to construct novel conformally flat metrics admitting higher rank Killing tensors.
Let us recapitulate the main features of the construction. A potential of the original dynamical system determines the conformal factor of the metric, as well as it relates the coordinate time to the proper time. Apart from Cartesian coordinates and a temporal variable, an extra spatial coordinate is needed in order to reproduce the conventional nonrelativistic energy conservation condition within the relativistic framework. In general, the metric admits three Killing vector fields, which form Lie algebra of the Poincaré group in $1+1$ dimensions. None of them is covariantly constant and, hence, the space–time does not belong to the Kundt class. If a dynamical system at hand admits an integral of motion polynomial in momenta, it can be uplifted to a Killing tensor. The uplift is different from that adopted within the conventional Eisenhart framework and essentially relies upon the condition that the geodesic is timelike.
Turning to possible further developments, it would be interesting to use the Killing vector fields in (13) to build the energy momentum tensor along the lines in [15] and study solutions to the corresponding Einstein equation.
A generalization to the case of time–dependent potentials is worth exploring as well.
Acknowledgements
This work was supported by the RF Ministry of Science and Higher Education under the project FSWW-2026-0010.

### v2

@@
 Anton Galajinsky Tomsk Polytechnic University, 634050 Tomsk, Lenin Ave. 30, Russia
-Tomsk State University of Control Systems and Radioelectronics,
-Lenin ave. 40, 634050 Tomsk, Russia
 e-mail: galajin@tpu.ru
@@
 In particular, geometrization of a Lagrangian conservative mechanics amounts to embedding Newton’s equation into the null geodesics of the Eisenhart metric [1, 2]
-$L(x,\dot{x})dt^{2}+dtds,$
-where $L(x,\dot{x})$ is the Lagrangian density, $t$ is the temporal coordinate, and $s$ is an extra variable [1, 2].
-The latter is needed in order to reproduce the conventional nonrelativistic energy conservation condition within the relativistic framework.
+$\mathcal{L}(x,\dot{x})dt^{2}+dtds,$
+where $\mathcal{L}(x,\dot{x})$ is the Lagrangian density, $t$ is the temporal coordinate, and $s$ is an extra variable [1, 2].
+The latter is needed in order to reproduce the conventional nonrelativistic energy conservation condition within the general relativistic framework.
 At the level of equations of motion, $s$ links to the action variable sometimes adopted in classical mechanics.
@@
 A salient feature of Eisenhart’s metric is that it admits a covariantly constant null Killing vector field $\frac{\partial}{\partial s}$, meaning that the space–time belongs to the Kundt class, as it contains a geodesic null congruence with vanishing expansion, shear and vorticity.
-If a Lagrangian density explicitly depends on time, $L(t,x,\dot{x})$, the metric is referred to as the Brinkmann–type metric or a plane–fronted wave with parallel rays (the pppp–wave).
+If a Lagrangian density explicitly depends on time, $\mathcal{L}(t,x,\dot{x})$, the metric is referred to as the Brinkmann–type metric or a plane–fronted wave with parallel rays (the pppp–wave).
 As is well known, in Riemannian geometry symmetries of a metric are described by Killing vectors, which generate infinitesimal coordinate transformations in space–time, $y^{\prime A}=y^{A}+\xi^{A}(y)$.
@@
 The importance of the Killing vectors and tensors is hard to overestimate.
-They give a clue to establishing complete integrability of the geodesic equations formulated in various black hole spacetimes, as well as allow one to separate variables in the Hamilton–Jacobi, Klein–Gordon and Dirac equations in strong gravitational fields (for a review and further references see [3]).
+They give a clue to establishing complete integrability of the geodesic equations formulated in various black hole space–times, as well as allow one to separate variables in the Hamilton–Jacobi, Klein–Gordon and Dirac equations in strong gravitational fields (for a review and further references see [3]).
 Apart from the aesthetic appeal, the Eisenhart lift provides an efficient means of constructing novel space–times with hidden symmetries.
-By applying the method to various integrable models one can build a plethora of examples of such a kind [4][16].
+By applying the method to various integrable models, one can build a plethora of examples of such a kind [4][16].
 Some of them provide exact solutions to the Einstein equations [4, 11, 12, 15].
@@
 [17][20] and references therein).
-The goal of this work is to study a variant of the Eisenhart lift, which is inspired by the Bohlin transformation [21] (for a review see [22]) connecting the planar harmonic oscillator to the Kepler problem.
+The goal of this paper is to study a variant of the Eisenhart lift, which is inspired by the Bohlin transformation [21] (for a review see [22]) connecting the planar harmonic oscillator to the Kepler problem.
 Given a dynamical system with $d$ degrees of freedom, a simple conformally flat metric on a (d+2)(d+2)–dimensional space–time of the Lorentzian signature is proposed, timelike geodesics of which include the original Newton equation.
@@
 By this reason, a numerical value of $g$ entering the original potential $V(x)=gU(x)$ is of no great importance.
+In what follows, we refer to $U$ as the prepotential.
 the Newton equation
@@
 which also means that the scalar curvature vanishes, harmonic prepotentials uplift to Ricci–flat metrics, and no cosmological term is allowed.
-Finally, assuming the original dynamical system (1) admits a polynomial integral of motion
+Finally, assuming the original dynamical system (1) admits a polynomial integral of motion of degree $n$ in velocities
 $\mathcal{I}=\sum_{k=0}^{n}I_{i_{1}\dots i_{k}}(x)\frac{dx_{i_{1}}}{dt}\dots\frac{dx_{i_{k}}}{dt},\qquad\frac{d\mathcal{I}}{dt}=0,$ (10)
@@
 [7, 10].
-of rank $n$ just by multiplying $\mathcal{I}$ with ${\left(\frac{dt}{d\lambda}\right)}^{n}$ and representing the latter in the form
-$K_{A_{1}\dots A_{n}}(y)\frac{dy^{A_{1}}}{d\lambda}\dots\frac{dy^{A_{n}}}{d\lambda}.$
-The resulting expression will automatically obey $\nabla_{(A_{1}}K_{A_{2}\dots A_{n+1})}=0$.
+of rank $n$.
+Recall that the geodesic equations (2), (5) imply that the coordinate time $t$ and the affine parameter $\lambda$ are affinely related $\frac{dt}{d\lambda}=\alpha$.
+By multiplying the conserved charge $\mathcal{I}$ with the constant $\alpha^{n}={\left(\frac{dt}{d\lambda}\right)}^{n}$, where $n$ is the highest power of the velocity in (10), one can bring the resulting expression to the homogeneous form
+$K_{A_{1}\dots A_{n}}(y)\frac{dy^{A_{1}}}{d\lambda}\dots\frac{dy^{A_{n}}}{d\lambda},$
+where $y^{A}=(ct,s,x_{1},\dots,x_{d})$, from which the components of the Killing tensor obeying $\nabla_{(A_{1}}K_{A_{2}\dots A_{n+1})}=0$ are read off.
 As was mentioned in the Introduction, within the context of general relativity Killing tensors describe hidden symmetries of a space–time.
@@
 Because Newton’s equation (1) holds invariant under the time reversal transformation $t\to-t$, the corresponding first integrals involve monomials in momenta which are all of even/odd power.
-Substituting $p_{i}=\frac{dx_{i}}{dt}=\frac{U}{\alpha}\frac{dx_{i}}{d\tau}$ into the integrals of motion and multiplying, where necessary, with the unity $1=\frac{1}{c^{2}}g_{AB}\frac{dy^{A}}{d\tau}\frac{dy^{B}}{d\tau}$, one obtains a homogeneous function $K_{A_{1}\dots A_{n}}\frac{dy^{A_{1}}}{d\tau}\dots\frac{dy^{A_{n}}}{d\tau}$, from which components of the Killing tensor are read off.
+Substituting $p_{i}=\frac{dx_{i}}{dt}=\frac{U}{\alpha}\frac{dx_{i}}{d\tau}$ into an integral of motion of degree $n$ in momenta and multiplying, where necessary, with the unity $1={\left(\frac{1}{c^{2}}g_{AB}\frac{dy^{A}}{d\tau}\frac{dy^{B}}{d\tau}\right)}^{m}$, where $m$ is a natural number, one can turn the latter into a homogeneous function $K_{A_{1}\dots A_{n}}\frac{dy^{A_{1}}}{d\tau}\dots\frac{dy^{A_{n}}}{d\tau}$ of degree $n$ in the velocities $\frac{dy^{A}}{d\tau}$, from which components of the Killing tensor $K_{A_{1}\dots A_{n}}$ are read off.
+In contrast to the conventional Eisenhart lift, higher rank Killing tensors of the Bohlin–type metric will always involve tsts–components.
 The Bohlin lift thus can be used to build novel conformally flat metrics of the Lorentzian signature admitting higher rank Killing tensors. Examples of such a kind are given in the next section.
@@
 where $a$ is a dimensionless constant and $b_{i}$ are constants having the dimension of inverse length.
-When Bohlin uplifted, it results in a metric solving the vacuum Einstein equations in the presence of a cosmological constant [^5] As a matter of fact, (23) yields a metric defined on a patch of the (d+2)(d+2)–dimensional anti de Sitter space, because $R_{ABCD}=(b_{i}b_{i})\left(g_{AC}g_{BD}-g_{AD}g_{BC}\right)$.
+When Bohlin uplifted, it results in a metric solving the vacuum Einstein equations in the presence of a cosmological constant [^5] As a matter of fact, (23) yields a metric defined on a patch of the (d+2)(d+2)–dimensional anti de Sitter space–time, because $R_{ABCD}=(b_{i}b_{i})\left(g_{AC}g_{BD}-g_{AD}g_{BC}\right)$.
+Anti de Sitter and de Sitter space–times thus fit nicely into the framework.
 $R_{AB}-\frac{1}{2}Rg_{AB}=\Lambda g_{AB},$
@@
 Our second example also belongs to the class of scaling invariant metrics and builds upon the Calogero model [24], which describes a set of $d$ identical particles on the real line interacting via the inverse–square potential
-$V(x_{1},\dots,x_{d})=\frac{1}{2}\sum_{i\neq j}^{d}\frac{g^{2}}{{(x_{i}-x_{j})}^{2}},$ (24)
+$V(x_{1},\dots,x_{d})=\frac{1}{2}\sum_{i,j=1,i\neq j}^{d}\frac{g^{2}}{{(x_{i}-x_{j})}^{2}},$ (24)
 where $g$ is a coupling constant. The model is integrable and the corresponding first integrals can be obtained from the Lax matrix (no sum over repeated indices below)
@@
 For example, the first four members in the string read
-$I_{1}=\sum_{i=1}^{d}p_{i},\quad I_{2}=\frac{1}{2}\sum_{i=1}^{d}p_{i}^{2}+\sum_{i<j}^{d}\frac{g^{2}}{{(x_{i}-x_{j})}^{2}},\quad I_{3}=\frac{1}{3!}\left(\sum_{i=1}^{d}p_{i}^{3}+3g^{2}\sum_{i<j}^{d}\frac{p_{i}+p_{j}}{{(x_{i}-x_{j})}^{2}}\right),$
-$I_{4}=\frac{1}{4!}\left(\sum_{i=1}^{d}p_{i}^{4}+4g^{2}\sum_{i<j}^{d}\frac{p_{i}^{2}+p_{j}^{2}+p_{i}p_{j}}{{(x_{i}-x_{j})}^{2}}+2g^{4}\sum_{i<j}^{d}\frac{1}{{(x_{i}-x_{j})}^{4}}+\right$.
-$\qquad\qquad\left.+4g^{4}\sum_{i\neq j,i\neq k,j<k}^{d}\frac{1}{{(x_{i}-x_{j})}^{2}{(x_{i}-x_{k})}^{2}}\right).$ (25)
+$I_{1}=\sum_{i=1}^{d}p_{i},\qquad I_{2}=\frac{1}{2}\sum_{i=1}^{d}p_{i}^{2}+\sum_{i,j=1,i<j}^{d}\frac{g^{2}}{{(x_{i}-x_{j})}^{2}},$
+$I_{3}=\frac{1}{3!}\left(\sum_{i=1}^{d}p_{i}^{3}+3g^{2}\sum_{i,j=1,i<j}^{d}\frac{p_{i}+p_{j}}{{(x_{i}-x_{j})}^{2}}\right),$
+$I_{4}=\frac{1}{4!}\left(\sum_{i=1}^{d}p_{i}^{4}+4g^{2}\sum_{i,j=1,i<j}^{d}\frac{p_{i}^{2}+p_{j}^{2}+p_{i}p_{j}}{{(x_{i}-x_{j})}^{2}}+2g^{4}\sum_{i,j=1,i<j}^{d}\frac{1}{{(x_{i}-x_{j})}^{4}}+\right$.
+$\qquad\qquad\left.+4g^{4}\sum_{i,j,k=1,i\neq j,i\neq k,j<k}^{d}\frac{1}{{(x_{i}-x_{j})}^{2}{(x_{i}-x_{k})}^{2}}\right).$ (25)
 Let us focus on the four–body case and construct a six–dimensional Bohlin–type metric, which admits irreducible Killing tensors of rank 3 and 4.
@@
 The prepotential which specifies the metric in (12) reads
-$U(x_{1},\dots,x_{d})=\frac{1}{2}\sum_{i\neq j}^{4}\frac{L^{2}}{{(x_{i}-x_{j})}^{2}},$
+$U(x_{1},\dots,x_{4})=\frac{1}{2}\sum_{i,j,i\neq j}^{4}\frac{L^{2}}{{(x_{i}-x_{j})}^{2}},$
 where $L$ is a constant having the dimension of length, which links to the coupling constant $g$ in (24) via $g^{2}=\frac{L^{2}}{2\alpha^{2}}$.
@@
 Substituting $p_{i}=\frac{dx_{i}}{dt}=\frac{U}{\alpha}\frac{dx_{i}}{d\tau}$ into $I_{3}$ and $I_{4}$ in (S0.Ex9) and multiplying the terms involving $g^{2}$ by the unity $1=g_{AB}\frac{dy^{A}}{d\tau}\frac{dy^{B}}{d\tau}$ and those containing $g^{4}$ by $1={\left(g_{AB}\frac{dy^{A}}{d\tau}\frac{dy^{B}}{d\tau}\right)}^{2}$, so as to obtain expressions homogeneous in $\frac{dy^{A}}{d\tau}$, one finally obtains the third–rank Killing tensor (no sum over repeated indices unless otherwise stated)
-$K_{tsi}=\frac{1}{2}L^{2}U^{2}\sum_{a,b,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}},$
-$K_{ijp}=U^{3}\delta_{ip}\delta_{jp}-\frac{1}{2}L^{2}U^{2}\left(\delta_{ij}\sum_{a\neq p}^{4}\frac{1}{{(x_{a}-x_{p})}^{2}}+\delta_{jp}\sum_{a\neq i}^{4}\frac{1}{{(x_{a}-x_{i})}^{2}}+\delta_{pi}\sum_{a\neq j}^{4}\frac{1}{{(x_{a}-x_{j})}^{2}}\right),$
+$K_{tsi}=\frac{1}{2}L^{2}U^{2}\sum_{a=1,a\neq i}^{4}\frac{1}{{(x_{a}-x_{i})}^{2}},$
+$K_{ijp}=U^{3}\delta_{ip}\delta_{jp}-\frac{1}{2}L^{2}U^{2}\left(\delta_{ij}\sum_{a=1,a\neq p}^{4}\frac{1}{{(x_{a}-x_{p})}^{2}}+\delta_{jp}\sum_{a=1,a\neq i}^{4}\frac{1}{{(x_{a}-x_{i})}^{2}}\right$.
+$\left.\qquad\quad+\delta_{pi}\sum_{a=1,a\neq j}^{4}\frac{1}{{(x_{a}-x_{j})}^{2}}\right),$
 as well as the fourth–rank Killing tensor
-$K_{ttss}=\frac{1}{6}L^{4}U^{2}\left(\sum_{a,b,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right),$
-$K_{tsij}=-\frac{1}{12}L^{4}U^{2}\delta_{ij}\left(\sum_{a,b,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right)$
-$+\frac{1}{6}L^{2}U^{3}\sum_{a,b,a\neq b}^{4}\frac{\left(2\delta_{ai}\delta_{aj}+\delta_{ai}\delta_{bj}\right)}{{(x_{a}-x_{b})}^{2}},$
+$K_{ttss}=\frac{1}{6}L^{4}U^{2}\left(\sum_{a,b=1,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c=1,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right),$
+$K_{tsij}=-\frac{1}{12}L^{4}U^{2}\delta_{ij}\left(\sum_{a,b=1,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c=1,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right)$
+$\qquad\qquad+\frac{1}{6}L^{2}U^{3}\sum_{a,b=1,a\neq b}^{4}\frac{\left(2\delta_{ai}\delta_{aj}+\delta_{ai}\delta_{bj}\right)}{{(x_{a}-x_{b})}^{2}},$
 $K_{ijkp}=U^{4}\sum_{a=1}^{4}\delta_{ai}\delta_{aj}\delta_{ak}\delta_{ap}$
-$+\frac{1}{12}L^{4}U^{2}\left(\delta_{ij}\delta_{kp}+\delta_{ik}\delta_{pj}+\delta_{ip}\delta_{jk}\right)\left(\sum_{a,b,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right)$
-$-\frac{1}{6}L^{2}U^{3}\left(2\sum_{a,b,a\neq b}^{4}\frac{\left(\delta_{ai}\delta_{aj}\delta_{kp}+\delta_{ai}\delta_{ak}\delta_{pj}+\delta_{ai}\delta_{ap}\delta_{jk}+\delta_{aj}\delta_{ak}\delta_{pi}+\delta_{aj}\delta_{ap}\delta_{ik}+\delta_{ak}\delta_{ap}\delta_{ij}\right)}{{(x_{a}-x_{b})}^{2}}\right$.
-$\left.+\sum_{a,b,a\neq b}^{4}\frac{\left(\delta_{ai}\delta_{bj}\delta_{kp}+\delta_{ai}\delta_{bk}\delta_{pj}+\delta_{ai}\delta_{bp}\delta_{jk}+\delta_{aj}\delta_{bk}\delta_{pi}+\delta_{aj}\delta_{bp}\delta_{ik}+\delta_{ak}\delta_{bp}\delta_{ij}\right)}{{(x_{a}-x_{b})}^{2}}\right).$
+$+\frac{1}{12}L^{4}U^{2}\left(\delta_{ij}\delta_{kp}+\delta_{ik}\delta_{pj}+\delta_{ip}\delta_{jk}\right)\left(\sum_{a,b=1,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c=1,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right)$
+$-\frac{1}{6}L^{2}U^{3}\left(2\sum_{a,b=1,a\neq b}^{4}\frac{\left(\delta_{ai}\delta_{aj}\delta_{kp}+\delta_{ai}\delta_{ak}\delta_{pj}+\delta_{ai}\delta_{ap}\delta_{jk}+\delta_{aj}\delta_{ak}\delta_{pi}+\delta_{aj}\delta_{ap}\delta_{ik}+\delta_{ak}\delta_{ap}\delta_{ij}\right)}{{(x_{a}-x_{b})}^{2}}\right$.
+$\left.+\sum_{a,b=1,a\neq b}^{4}\frac{\left(\delta_{ai}\delta_{bj}\delta_{kp}+\delta_{ai}\delta_{bk}\delta_{pj}+\delta_{ai}\delta_{bp}\delta_{jk}+\delta_{aj}\delta_{bk}\delta_{pi}+\delta_{aj}\delta_{bp}\delta_{ik}+\delta_{ak}\delta_{bp}\delta_{ij}\right)}{{(x_{a}-x_{b})}^{2}}\right).$
+It is straightforward to verify that the Killing tensors are irreducible, i.e. they do not decompose into a combination of the Killing vector fields (13) and the metric (12).
 In a similar fashion, by increasing the number of particles in the original Calogero model, one can build a (d+2)(d+2)–dimensional Bohlin–type metric admitting Killing tensors of rank up to $d$.
 5. Conclusion
-To summarize, inspired by the Bohlin transformation relating the planar harmonic oscillator to the Kepler problem, a variant of the Eisenhart lift was proposed, in which a Lagrangian conservative dynamical system with $d$ degrees of freedom was embedded into timelike geodesics of a conformally flat metric defined on a (d+2)(d+2)–dimensional space–time of the Lorentzian signature.
+To summarize, inspired by the Bohlin transformation relating the planar harmonic oscillator to the Kepler problem, a variant of the Eisenhart lift was proposed, in which a Lagrangian conservative dynamical system with $d$ degrees of freedom was embedded into timelike geodesics of a conformally flat stationary metric defined on a (d+2)(d+2)–dimensional space–time of the Lorentzian signature.
 The (anti) de Sitter space–time fitted nicely into the framework.
@@
 A potential of the original dynamical system determines the conformal factor of the metric, as well as it relates the coordinate time to the proper time.
-Apart from Cartesian coordinates and a temporal variable, an extra spatial coordinate is needed in order to reproduce the conventional nonrelativistic energy conservation condition within the relativistic framework.
-In general, the metric admits three Killing vector fields, which form Lie algebra of the Poincaré group in $1+1$ dimensions.
+Apart from Cartesian coordinates and a temporal variable, an extra spatial coordinate is needed in order to reproduce the conventional nonrelativistic energy conservation condition within the general relativistic framework.
+In general, the metric admits three Killing vector fields, which form the Lie algebra of the Poincaré group in $1+1$ dimensions.
 None of them is covariantly constant and, hence, the space–time does not belong to the Kundt class.
@@
 Acknowledgements
-This work was supported by the RF Ministry of Science and Higher Education under the project FSWW-2026-0010.
+This work was supported by the RF Ministry of Science and Higher Education under the project FSWW-2026-0046.